OpenStudy (anonymous):

Can someone tell me if I solved, 2x^2 + 13x + 15 correctly? The first trinomial with be 2x^2 + 13x + 15. To solve you first you check to see if there is a GCF. For this one there is none. Next you are going to multiply the first coefficient which is 2 with the last term of the trinomial which is 15. So 2 * 15 = 30. Next you need to find factors of 30 that add to give you the coefficient of the middle term which is 13. I used -2 + 15. So now it is, 2x^2 - 2x + 15x + 15. Now I need to separate it into two groups giving me, 2x^2 - 2x and 15x + 15. Now I need to find the GCF of the first gro

OpenStudy (anonymous):

group which is 2x. And the GCF of the second group is 15. So now it is 2x(x - 1) 15(x + 1). Finally I factor the common binomial which is (x - 1)(2x + 15).

OpenStudy (anonymous):

it is wrong..

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

Your question is wrong I think.. There will come -13x..

OpenStudy (cwrw238):

the factors you have found expand to 2x^2 + 13x - 15 (not + 15)

OpenStudy (anonymous):

Check your question once again..

OpenStudy (anonymous):

For \(2x^2 - 13 + 15\), factors will be -10 and -3.. \[\large 2x^2 - 10x - 3x +15 \rightarrow2x(x-5) -3(x-5) \rightarrow (2x-3)(x-5)\] For \(2x^2 + 13x - 15\), factors will be +15 and -2, \[\large 2x^2 - 2x + 15x -15 \rightarrow 2x(x-1) + 15(x-1) \rightarrow (x-1)(2x+15)\]

OpenStudy (anonymous):

\[(2x ^{2}+13x+15)=2x ^{2}+10x+3x+15\] =2x(x+5)+3(x+5) =(2x+3)(x+5)

OpenStudy (anonymous):

Aadmi jab jyada pad leta hai to uska dimaag uske ghutnon me chale jaat hai @nitz Tumhe nahin keh raha hu apne aap ko keh raha hu..

OpenStudy (anonymous):

whatever@waterineyes